1. Two samples of a radioisotope were spilled in a nuclear laboratory. The activity of one sample was 8 kBq and that of the other was 17 mCi .

Convert 17 mCi to kilobecquerels so that you can directly compare the activity of each sample.

Express your answer using two significant figures.
Answer #1

1 Ci = 3.7\small \times 1010 Bq

1mCi = 10-3 Ci

1 Bq = 10-3 KBq

so,

17 mCi = 17 (mci) \small \times10-3(ci) 1(mCi)  \small \times  3.7 x 1010(Bq) 1(Ci)\small \times10-3(K Bq) 1(B )

= 17 \small \times 10-3\small \times 3.7 \small \times 1010\small \times 10-3 KBq

= 629000 KBq

= 6.3 \small \times105 KBq ( 2 significant figure)